wbact is the AHeuap) of c liqu Pressure of le torr at 82.3ee and a boiling oint of 93. at Latm
Answer #1

as we know that 1 atm = 760 torr   

H is enthalpy of vaporisation

R =  8.314472 J mol1 K1

given T1 = 97.2 + 273 K = 370.2 K

T2 = 82.3+273K = 355.3 K

P2 = 617 torr

P1 = 760 torr

therefore

using clausius clapeyron equation
ln[ P2 / P1 ] = [ H / R ][ (1 / T1) - (1 / T2) ]

ln [617/760] = [H/8.314472][(1/370.2) - (1/355.3)]

therefore H = 17290 J/mol